OpenStudy (anonymous):

Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form. y – 2 = 3(x – 5) y + 8 = 3(x + 5) y + 5 = 3(x + 2) y – 5 = 3(x – 2)

OpenStudy (fibonaccichick666):

First, can you tell me what is point-slope form?

OpenStudy (anonymous):

y=mx+b

OpenStudy (fibonaccichick666):

That is the standard eq. of a line, but not point-slope form any other ideas?

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

???

OpenStudy (anonymous):

FibonacciChick666

OpenStudy (precal):

you have to put the @ symbol in front of the user name to send the notification

OpenStudy (precal):

you know that m=3 because you were told slope is 3

OpenStudy (precal):

(2,5) implies x=2 and y =5 so just sub it into the second formula you stated

OpenStudy (anonymous):

so wats the answer? @precal

OpenStudy (precal):



OpenStudy (precal):

just sub those values

OpenStudy (anonymous):

is it c?

OpenStudy (precal):